OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: y less than or equal to –x + 7 y < 3x + 11 Answer (–5, 5) (5, 5) none of the above (–5, –5)

OpenStudy (anonymous):

what is the answer

OpenStudy (ash2326):

We have \[y \le -x+7\ and\ y<3x+11\] Let's check each option. The point should simultaneously satisfy both a)(-5, 5) x=-5 y=5 \[5 \le -(-5)+7\ and\ 5<3*(-5)+11\] First is satisfied but second isn't Let's check c)(5, 5) \[5 \le -5+7\ and\ 5<3*5+11\] here second is satisfied but first isn't Let's check d d)(-5, -5) \[-5 \le -(-5)+7\ and\ -5<3*-5+11\] Here both are satisfied so option D is the correct answer